a garrison of 1600 men is provisioned for 18 weeks at the rate of 25 kg per day per man.if some more persons join the garrison, the provisions would last for 16 weeks at the rate of 20 kg per day per man. find the number of persons joining the garrison.
Found 2 solutions by mananth, robertb:
Answer by mananth(16949)

m1=1600: w1=18: k1=25
...
...
m2=x: w2=16: k2=20
...
m2= 1600*(18/16)*(25/20)
m2= 2250
..
2250-1600 = 650 men joined the garrison
...
m.ananth@hotmail.ca

Answer by robertb(5830)

1600*25*126 = (1600 + x)20*112
2250 = 1600+x;
x = 650, the number of people joining the garrison.